Given one point on the graph (0, -11) and the slope, m = -2, develop an equation in the slope-intercept format.

Answer by Alan3354(69443)

A line and a point example.
-------------------
The slope, m = -3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note, or at Moral Loophole@aol.com